OpenStudy (anonymous):

The sum of the digits of a two-digit number is 14. If the number represented by reversing the digits is subtracted from the original number, the result is 18. What is the original number?

OpenStudy (amistre64):

x+y = 14 (10x+y) - (10y+x) = 18

OpenStudy (anonymous):

Ah I see, I was writing it completely wrong

OpenStudy (amistre64):

or .. (7+x)+(7-x) = 14 10(7+x)+(7-x) -10(7-x)-(7+x) = 18 :)

OpenStudy (anonymous):

9x - 9y = 18?

OpenStudy (amistre64):

if you wanna go that rout, then that good so far

OpenStudy (amistre64):

multiply the first eq by 9 and add them together

OpenStudy (amistre64):

or divide out the 9 from that .... either way

OpenStudy (anonymous):

thanks, i got the correct answer

OpenStudy (amistre64):

yay!! ... was it 86?

OpenStudy (anonymous):

yep!